# Cross-Problem Solving for Network Optimization: Is Problem-Aware Learning the Key?

### v1

@@ V-A2 Model Settings
For the encoder-decoder, we use a series of fully connected layers to embed the input parameter values into a hidden_dim-dimensional space, and an nn.Embedding layer to embed the parameter types into the latent space. During decoding, each token is decoded using several fully connected layers. The parameter values are predicted with a Sigmoid activation, while the parameter types are predicted with a Softmax activation. The encoder-decoder was trained following Algorithm 1 for 20 epochs with a batch size 256. The initial learning rate was set to 0.00015 and adjusted using a cosine annealing schedule.
For the PAD backbone, the Transformer is configured with a hidden dimension of 128, 8 attention heads, 6 encoder layers and 6 decoder layers. In the subsequent experiments, the PAD backbone is trained following Algorithm 3 for 50 epochs with a batch size of 2048. The diffusion setting adopts a Gaussian diffusion mechanism with 50 diffusion steps, cosine noise scheduling, DDIM [51] acceleration with 5 generation steps, and 2 parallel sampling runs during inference. In practice, many existing solution generation methods require tool-integrated reasoning or decoding strategies to further interpret the decoded solutions [22, 52, 34]. Our approach introduces a simple outside decoding step based on the absolute magnitude of each generated solution element prior to de-normalization, without relying on any external reasoning tools.
For the constraint-aware module, it is implemented using only a TransformerEncoder architecture with the same parameter settings as the PAD backbone. In the experiments, it is trained according to Algorithm 2 for 10 epochs, with an additional 2-epoch fine-tuning applied to a few more complex problems. The batch size is set to 1024. Since uniformly sampled solutions for some problems may exhibit a strong imbalance between valid and invalid samples (e.g., over 80% of the training samples are labeled as violating constraints), which harms the model’s generalization ability. We apply oversampling during training to balance the distribution. Additionally, we assign different weights to the two classes in the CrossEntropyLoss, with a weight of 1.1 for label 0 and 1.0 for label 1. No oversampling is used during testing.
In training the backbone with the inclusion of constraint loss, two variables are considered: the method of computing the constraint loss and the epoch at which the constraint loss is introduced. We design two strategies for computing the constraint loss. The first, referred to as “default”, directly uses the output of the constraint-aware module as the loss value. The second, termed “t_increase”, applies a linear scaling to the module’s output based on the denoising timestep, assigning larger losses to noisy solutions closer to timestep 0. For each loss computation method, we explore five different epoch ratios for introducing the constraint loss: [0.2, 0.4, 0.6, 0.8, 1.0]. Each value denotes the proportion of the total training epochs during which the constraint loss is applied at the end of training. For example, the value 0.2 means that the constraint loss is applied only during the final 10 epochs of a 50-epoch training process. In total, this brings 10 distinct settings for constraint loss integration.
@@ V-A3 Metrics
  • GT_GAP: This metric directly measures solution quality by computing the absolute difference between 1 and the ratio of the generated solution’s objective value to the ground-truth optimal value, i.e., $\mathrm{GT\_GAP}=|\frac{f(\mathbf{x},\tilde{\mathbf{y}})}{f(\mathbf{x},\mathbf{y})}-1|$. The final score is the average GT_GAP across all samples in the test set.
  • AVG_GT_GAP: For more concise reporting, this simplified version of GT_GAP directly refers to the average GT_GAP across all test problems.
@@
  • CONS_IF: This metric evaluates whether at least one valid solution exists among the parallel samples for each input. A value of 1 is assigned if any valid solution exists, and 0 otherwise. The final score is the average over all test inputs.
@@
V-B Analysis of Ranked Pool
@@ V-B1 Specific Settings
To empirically validate the effectiveness of the proposed ranked pool method for extracting fixed-length problem embeddings, we conduct a controlled comparison against two widely used pooling strategies: MaxPool and AvgPool. All evaluations are performed under a consistent experimental setup trained on (P2,P3,P5,P6), with the constraint-aware module disabled to isolate the effects of the token pooling mechanism.
@@ V-B2 Results
In terms of solution quality, measured by the average GT_GAP across all test problems, the ranked pool shows consistent improvements over the baseline methods. Specifically, ranked pool yields an AVG_GT_GAP of 0.1263, compared to 0.1350 for MaxPool and 0.1370 for AvgPool. It suggests that the tokens selected via ranked pool better preserve the symbolic and structural nuances of the optimization problems, which are essential for accurate solution generation.
Figure 5: Comparison on GT_GAP between models with and without problem feature embedding from 10 different training problem selections. Each line plot includes a light-colored shaded band representing the error bar given by the 95% confidence interval of the corresponding metric across all test samples. The line itself indicates the mean value. In the two rows of subplots, the training problem selection strategy for the first row is “ascending,” while for the second row it is “descending.” The specific training problems used for each subplot are indicated in the corresponding titles.
Figure 6: Comparison on CONS_ACC between models with and without problem feature embedding from 10 different training problem selections. Each line plot includes a light-colored shaded band representing the error bar given by the 95% confidence interval of the corresponding metric across all test samples. The line itself indicates the mean value. In the two rows of subplots, the training problem selection strategy for the first row is “ascending,” while for the second row it is “descending.” The specific training problems used for each subplot are indicated in the corresponding titles.
V-B3 Analysis of Problem Distinguishability
To better understand the source of the observed performance differences, we analyze the ability of each pooling method to generate problem embeddings that distinguish between different optimization problems. We hypothesize that ranked pool promotes discriminative representations by preserving structural and symbolic distinctions across problems.
To verify this, we compute pairwise cosine similarities between the pooled MathBERT embeddings of all 6 problems with different pool methods. For each pool method—AvgPool, MaxPool, and Ranked Pool—we apply the respective method to the original embeddings and calculate the resulting similarity matrix. The heatmaps are shown in Fig. 4.
The results reveal a clear pattern. MaxPool produces extremely high similarity scores (all above 0.98), indicating overly homogeneous embeddings that fail to differentiate between problems. AvgPool shows slightly more variation, but the similarities remain uniformly high, with most values above 0.90. In contrast, the ranked pool similarity matrix exhibits much greater variability: while semantically related problems (e.g., P1 and P4) retain moderately high similarity (above 0.5), unrelated problems show significantly lower scores, some as low as 0.1–0.3. This confirms that ranked pool can better reflect symbolic and structural problem differences.
V-B4 Summary
These findings support our hypothesis that token-level similarity analysis provides a more principled approach to selecting semantically rich and diverse tokens, especially in mathematical expressions where subtle symbolic differences matter. Unlike MaxPool, which may overemphasize frequent or high-activation tokens, and AvgPool, which tends to oversmooth, the ranked pool explicitly promotes token diversity and structural sensitivity through similarity-based ranking.
Overall, the ranked pool method offers a lightweight yet effective strategy for compressing long, symbolic representations into fixed-length embeddings. Its advantage lies not only in better capturing relevant token information but also in enhancing the distinguishability between problem instances for improving generalization in mathematical optimization tasks.
Figure 7: Comparison on GT_GAP between models before and after incorporating constraint loss from 10 different training problem selections. Each line plot includes a light-colored shaded band representing the error bar given by the 95% confidence interval of the corresponding metric across all test samples. The line itself indicates the mean value. All subplots share a common y-axis scale. In the two rows of subplots, the training problem selection strategy for the first row is “ascending,” while for the second row it is “descending.” The specific training problems used for each subplot are indicated in the corresponding titles.
Figure 8: Comparison on CONS_ACC between models before and after incorporating constraint loss from 10 different training problem selections. Each line plot includes a light-colored shaded band representing the error bar given by the 95% confidence interval of the corresponding metric across all test samples. The line itself indicates the mean value. In the two rows of subplots, the training problem selection strategy for the first row is “ascending,” while for the second row it is “descending.” The specific training problems used for each subplot are indicated in the corresponding titles.
@@ V-C1 Specific Settings
To illustrate the effectiveness of incorporating problem feature embedding, we conduct experiments under the default training setup, exploring two variables: (1) whether problem feature embedding is included, and (2) the selection of the training problem set. This results in a total of 20 different experimental settings. We then evaluate the model using two metrics: GT_GAP and CONS_ACC.
V-C2 Results
As shown in Fig. 5 and Fig. 6, models incorporating problem embedding (blue lines) generally outperform those without it (red lines), except in the extreme case where only a single problem is used for training. It can be observed that models without problem embedding exhibit higher variance on both metrics, particularly on GT_GAP. The model’s performance on unseen problems in the training set further reflects its cross-problem generalization capability. For models without problem embedding, any degree of generalization is largely attributable to the benefits of parallel sampling and outside decoding. However, these models perform noticeably worse on CONS_ACC, since without problem embedding, the model lacks even the basic ability to identify the problem, let alone learn its characteristics.
The training problem selection for each subplot in Fig. 5 and Fig. 6 follows a specific strategy: the first row of subplots corresponds to the “ascending” strategy, while the second row corresponds to the “descending” strategy. Regarding the two problem selection strategies, the results show that “descending” consistently performs better. We believe the reason is that “descending” selects problems with the highest total similarity scores, effectively choosing the most representative and diverse problems from the problem space. This also suggests that the higher the similarity between problems, the stronger the model’s potential for cross-problem generalization.
Regarding the number of training problems, it is naturally observed that more training problems lead to better overall performance. A few small exceptions exist-for instance, in Fig. 5, the two subplots “Trained on (P1,P4,P5,P6)” and “Trained on (P1,P4,P5)” show that for test problem index 12, the model trained on four problems slightly underperforms the one trained on three. This is mainly due to conflicting gradient directions among different problems during training; when more problems are jointly trained, performance trade-offs may arise. Although some multi-task learning techniques exist to mitigate such gradient conflicts, addressing this is beyond the scope of our current study.
Regarding problem characteristics, some of the six problems are inherently easier or more difficult. For instance, P1 is a constrained convex optimization problem, making it relatively simple, thus all models perform well on it. In contrast, P4 is a constrained integer programming problem, which is more complex. Combined with convergence issues caused by gradient conflicts across different problems, this leads to generally poor performance on P4 for all models.
As for cross-scale generalization, we observe a general trend where performance deteriorates as the problem dimension increases. This may be due to the data’s resource configuration. While the total amount of network resources remains nearly constant, the number of optimization variables increases, and the constraints remain unchanged. This significantly shrinks the feasible region and increases the problem’s complexity, impacting model performance.
V-C3 Summary
@@
Overall, the results in Fig. 5 and Fig. 6 demonstrate that incorporating problem feature embedding significantly enhances the model’s ability to generalize across different problems. This improvement is evident in both the quality and feasibility of the generated solutions, highlighting the strong potential of problem-aware learning.
@@ V-D1 Specific Settings
To demonstrate the effectiveness of the constraint-aware module, we compare the PAD model’s performance before and after introducing the constraint loss. Through repeated experiments, we identify the best constraint loss configuration using the “default” loss computation method with an epoch introduction ratio of 0.6. We fix this configuration and conduct repeated experiments under 10 different training problem selection settings, comparing performance on the GT_GAP and CONS_ACC metrics. Additionally, we illustrate the constraint satisfaction classification performance of the shared frozen constraint-aware module using three metrics: ACC, TPR and TNR.
@@ V-D2 Results
As shown in Table LABEL:tab_cons_cls_performance, the constraint-aware module achieves consistently high cross-problem classification accuracy (mostly above 0.95), particularly in identifying valid solutions, with TNR approaching 1 across the board. The classification accuracy is slightly lower for a few problems, such as P3, which only involves simple constraints like ensuring that allocated resources do not exceed the total available. We attribute this drop primarily to conflicting gradient directions during multi-problem training. Nevertheless, the overall accuracy of the constraint-aware module is sufficiently high for use in constraint loss computation, which effectively penalizes invalid solutions and rarely misclassifies valid ones. After introducing the constraint loss, CONS_IF improved from mostly above 0.8—with some cases as low as 0.64—to consistently above 0.92. With a slight increase in the number of parallel samples (e.g., to 4), it can approach 1.
As shown in Fig. 7 and Fig. 8, models incorporating the constraint loss (blue lines) generally perform better, particularly in terms of CONS_ACC, where improvements of up to 60% are observed for certain problems. The increase in CONS_ACC allows the parallel sampling process to cover a broader space of valid solutions, making it easier to reach better overall solutions and thus directly improving GT_GAP as well.
In some cases, the improvement in CONS_ACC plateaus after introducing the constraint loss. We believe this is due to the upper performance bound imposed by joint training across multiple problems. Additionally, in certain cases, the change in CONS_ACC is minimal while GT_GAP improves significantly, such as P4 in the “Trained on (P2, P3)” subplots. This suggests that the constraint-aware module’s implicit understanding of problem-specific constraints may also contribute to the quality of the generated feasible solutions.
TABLE I: The binary classification performance of the constraint-aware module.

### v2

@@ V-A2 Model Settings
Table I presents the main hyperparameter settings of PAD. Unless otherwise specified, all experimental figures and tables are based on these configurations. The settings of diffusion-related parameters such as diffusion_steps and inference_diffusion_steps in Table I are derived from existing engineering analyses on the impact of hyperparameters in diffusion-based solution generation [60, 34]. For a more detailed understanding, this subsection provides additional descriptions of the encoder–decoder, the constraint-aware module, and the latent diffusion setup.
For the encoder–decoder, a series of multilayer perceptrons are used to map input parameter values into a high-dimensional latent space (dimension denoted as hidden_dim) or decode from it. The entries encode_dims and decode_dims in Table I describe the progressive encoding and decoding dimensions, from 1D parameter values to the hidden_dim latent representation. Integer-type input parameters are handled via an nn.Embedding layer. Across the ten problems, there are 19 distinct parameter and variable types with different physical meanings, as detailed in our open-source repository. The parameter values are predicted using a Sigmoid activation, while parameter types are predicted using a Softmax activation. Due to space limitations and the auxiliary role of the encoder–decoder, we did not dedicate a separate section to its performance results. Here, we only note that its element-value MSE loss during testing is approximately $10^{-7}$, and the prediction accuracy of physical types is close to 1.0.
For the constraint-aware module, we employ a TransformerEncoder backbone. Since uniformly sampled solutions for certain problems may exhibit severe imbalance between valid and invalid samples (e.g., over 80% of the training samples violate constraints), which harms generalization, we apply oversampling during training to balance the distribution. Moreover, after each training epoch, we update the cross-entropy loss weights based on the model’s performance in the previous epoch: the class of 0-1 with poorer performance is assigned a higher weight (up to 1.5), while the better-performing class is assigned a lower weight (down to 1).
For the PAD backbone, we employ a full Transformer architecture. Diffusion noise is scheduled using a cosine scheduler, and denoising diffusion implicit models (DDIM) [61] acceleration is applied during sampling. In practice, many existing solution generation methods require tool-integrated reasoning or decoding strategies to further interpret the decoded solutions [23, 60, 35]. Our approach introduces a simple outside decoding step based on the absolute magnitude of each generated solution element prior to de-normalization, without relying on any external reasoning tools. Constraint losses are introduced only during the final epochs of training; the hyperparameter constrain_loss_start_ratio in Table I specifies the proportion of the training process during which the constraint loss is applied. For example, (constrain_loss_start_ratio=0.5) indicates that the constraint loss is incorporated throughout the latter half of training. This setting is empirical: introducing the loss too early may cause gradient competition that hinders convergence of the diffusion loss, while introducing it too late may result in undertraining.
The batch sizes for the encoder–decoder, constraint-aware module, and PAD’s latent diffusion backbone are 256, 128, and 352, respectively. These settings are primarily determined by hardware limitations (notably GPU memory) and aim to maximize training efficiency, so they are not included in Table I due to their limited generalizability. The constraint-aware module uses the smallest batch size because oversampling doubles the number of samples per problem dimension relative to the original dataset, thereby consuming more GPU memory. For learning rates, we employ a cosine annealing schedule across all models. In addition, to mitigate gradient conflicts between different problems, we aggregate losses across all training problems in a single batch and perform backpropagation jointly, rather than using gradient descent on single-problem batches.
For the parallel sampling ($p_{\rm num}$) setup during inference, the value of 2 in Table I is chosen to amplify the impact of different experimental variables on solution quality. If the parallel sampling count were increased, the performance differences in the figures would become less pronounced, making them harder to observe and analyze.
@@ V-A3 Metrics
  • GT_GAP: This metric directly measures solution quality by computing the absolute difference between 1 and the ratio of the generated solution’s objective value to the ground-truth optimal value, i.e., $\mathrm{GT\_GAP}=|\frac{f(\mathbf{x},\tilde{\mathbf{y}})}{f(\mathbf{x},\mathbf{y})}-1|$. The final score is the average GT_GAP across all samples in the test set of one problem.
  • GLOB_GT_GAP: For more concise reporting, this simplified version of GT_GAP directly refers to the average GT_GAP across all test problems.
@@
  • CONS_IF: This metric evaluates whether at least one valid solution exists among the parallel samples for each input. A value of 1 is assigned if any valid solution exists, and 0 otherwise. The final score is the average over all test inputs of one problem.
Figure 4: Comparison of different pooling methods on GT_GAP $\downarrow$ and CONS_IF $\uparrow$. The three subplots correspond to the three pooling lengths. Yellow markers represent GT_GAP (measured on the left y-axis), while blue markers represent CONS_IF (measured on the right y-axis).
@@
TABLE II: GLOB_GT_GAP $\downarrow$ performance across different pooling methods and pooling lengths ($m$). Here, “Total” denotes the test results aggregated over all problems, “Train” refers to the test results on training problems only, and “Test” indicates the results on problems outside the training set.
Rank Pool Max Pool Avg Pool
Total, $m=12$ $9.58\times 10^{-2}$ $8.00\times 10^{-2}$ $9.58\times 10^{-2}$
Total, $m=14$ $8.00\times 10^{-2}$ $1.08\times 10^{-2} 11$$1.06\times 10^{-2}$
Total, $m=20$ $\mathbf{7.64\times 10^{-2}}$ $8.98\times 10^{-2}$ $7.73\times 10^{-2}$
Train, $m=12$ $8.85\times 10^{-2}$ $9.00\times 10^{-2}$ $10.29\times 10^{-2}$
Train, $m=16$ $8.62\times 10^{-2}$ $13.06\times 10^{-2}$ $14.65\times 10^{-2}$
Train, $m=20$ $8.45\times 10^{-2}$ $10.58\times 10^{-2}$ $\mathbf{8.13\times 10^{-2}}$
Test, $m=12$ $10.07\times 10^{-2}$ $7.33\times 10^{-2}$ $9.11\times 10^{-2}$
Test, $m=16$ $7.59\times 10^{-2}$ $9.75\times 10^{-2}$ $8.66\times 10^{-2}$
Test, $m=20$ $\mathbf{7.11\times 10^{-2}}$ $7.92\times 10^{-2}$ $7.46\times 10^{-2}$
V-B Analysis of Rank Pool
@@ V-B1 Specific Settings
To empirically validate the effectiveness of the proposed rank pool method for extracting fixed-length problem embeddings, we conduct experiments using three different pooling methods—Rank Pool, MaxPool, and AvgPool—to directly compare their impact on solution quality. In addition, we design three groups of experiments with different pooling lengths ($m=12,16,20$) to analyze how it affects solution quality. Except for variations in the pooling method and pooling length, all models are constructed according to the settings in Table I, with the constraint-aware module introduced under the same configuration.
Figure 5: Comparison of different pooling lengths on GT_GAP $\downarrow$ and CONS_IF $\uparrow$. The three subplots correspond to the three pooling methods. Yellow markers represent GT_GAP (measured on the left y-axis), while blue markers represent CONS_IF (measured on the right y-axis).
Figure 6: Comparison of with and without problem embeddings and constraint loss on GT_GAP $\downarrow$ and CONS_IF $\uparrow$.
@@ V-B2 Results
The comparative results for pooling methods and pooling lengths are presented in Table LABEL:tab_total_gt_gap, Fig. 4 and 5.
Table LABEL:tab_total_gt_gap highlights the clearest comparison between pooling methods and lengths. Regarding pooling length, both rank pool and AvgPool exhibit a trend where larger $m$ consistently improves performance. In contrast, MaxPool behaves unusually. The MaxPool performs worst at $m=16$ but best at $m=12$, which may be attributed to the embedding characteristics of the selected problems, where fragmenting tokens through pooling yields some distinguishable representations at smaller $m$. In terms of pooling methods, rank pool achieves the best performance on global problems (“Total”) and unseen problems (“Test”) as $m$ increases, but is slightly outperformed by AvgPool on training problems (“Train”). This may be because once $m$ becomes sufficiently large, information is already well preserved, reducing the negative learning effect of AvgPool’s smoothing operation and allowing it to achieve a better balance.
@@ V-C1 Specific Settings
To illustrate the effectiveness of incorporating problem feature embedding, we conduct experiments under the default training setup of Table I, exploring the impact of including or excluding problem feature embedding on performance.
Fig. 4 and Fig. 5 present dual-metric comparisons of GT_GAP and CONS_IF across different problems with respect to pooling methods and pooling lengths, respectively. However, due to variations in problem difficulty and the randomness introduced when $p_{\rm num}=2$, the overall advantages are not immediately obvious, and these figures serve better as references for observing the overall performance range. In Fig. 4, the Rank pool is seen to achieve the best results on both GT_GAP and CONS_IF for most problems. Fig. 5 shows a trend largely consistent with Table I, where only the rank pool exhibits a tendency for performance to improve as $m$ increases.
V-C2 Summary
These findings highlight the effectiveness of pooling original problem feature embeddings into a fixed length under limited model size and computational resources, providing empirical guidance for the choice of pooling methods and pooling lengths.
Overall, the rank pool method offers a lightweight yet effective strategy for compressing long, symbolic representations into fixed-length embeddings. Its advantages lie not only in capturing critical token-level information but also in offering better support for the model’s final solution generation, particularly demonstrating stronger performance on unseen problems outside the training set.
V-C3 Results
As shown in Fig. 6, on GT_GAP, the model with problem embedding (yellow triangles) outperforms the one without problem embedding (yellow crosses) on most problems, particularly demonstrating better generalization on out-of-training problems (P5 and P8). On P6, P9, and P10, however, the results without problem embedding are slightly better. We attribute this to the limited model capacity, which amplifies gradient competition across different problems, thereby increasing the learning difficulty when problem embeddings are included and leading to weaker performance in certain cases. In conjunction with the findings of Section IV-A, we conjecture that P6, P9, and P10 may inherently allow the problem structure to be inferred directly from the input parameters. As a result, excluding problem embedding reduces training complexity while having only a minor impact on problem identification. For CONS_IF, neither the model with problem embedding (blue triangles) nor the one without it (blue crosses) incorporates constraint loss. Nonetheless, the explicit mathematical structure and problem-discriminative information introduced by problem embeddings improve the feasibility rate of generated solutions on some problems.
V-C4 Summary
@@
Overall, the results in Fig. 6 demonstrate that incorporating problem feature embedding effectively enhances the model’s ability to generalize across different problems. This improvement is evident in both the quality and feasibility of the generated solutions, highlighting the potential of problem-aware learning.
@@ V-D Performance of Constraint-Aware Module
Figure 7: Performance of the constraint-aware module in terms of ACC $\uparrow$, TPR $\uparrow$, and TNR $\uparrow$ on training problems.
Figure 8: Comparison of with and without the constraint-gated training mechanism on GT_GAP $\downarrow$ and CONS_IF $\uparrow$.
@@ V-D1 Specific Settings
To demonstrate the effectiveness of the constraint-aware module, we compare the PAD model’s performance with and without introducing the constraint loss. Furthermore, we conduct ablation studies on the constraint-gated loss mechanism across all three pooling strategies to assess its effectiveness. Additionally, we illustrate the constraint satisfaction classification performance of the shared frozen constraint-aware module using three metrics: ACC, TPR, and TNR.
@@ V-D2 Results
As shown in Fig. 6, when comparing models trained with constraint loss (yellow circles) and without constraint loss (yellow triangles), we observe a clear GT_GAP improvement across almost all problems, with only a few cases showing no gain, likely due to limited model capacity. This improvement primarily stems from the significantly higher proportion of feasible solutions generated after introducing constraint loss into the training process. This can also be observed from the comparison of CONS_IF under constraint loss (blue circles) and without it (blue triangles). The increased feasibility ratio directly enhances the efficiency of parallel sampling in exploring high-quality solution distributions, thereby yielding a stable improvement in overall solution quality.
As shown in Fig. 7, our constraint-aware module is used only during training, and thus we report its discrimination accuracy on the four training problems. The module achieves accuracy above 0.9 across all three metrics—ACC, TPR, and TNR. As a differentiable surrogate for the original non-differentiable constraint functions, such accuracy is already sufficient. High TPR and TNR help minimize the risk of penalizing feasible solutions or overlooking infeasible ones, thereby preventing the diffusion model from learning incorrectly.
As shown in Fig. 8, introducing the constraint-gated loss mechanism (yellow circles) yields only marginal GT_GAP improvement compared with not using it (yellow triangles), and for rank pool and AvgPool it even leads to opposite effects on certain problems (P9, P10). We suspect this is because half of the training resources originally allocated to full problem embedding and constraint loss are diverted to incomplete embeddings with constraint loss disabled, resulting in insufficient training under full constraints. However, the mechanism shows significant improvements on CONS_IF in certain problems, indicating its effectiveness.
Figure 9: Effect of parallel sampling $p_{num}$ on GT_GAP $\downarrow$ and CONS_IF $\uparrow$.
V-D3 Summary
In summary, the results presented in Fig. 6, Fig. 7 and 8 demonstrate that incorporating constraint loss effectively improves the feasibility of the solutions generated by the PAD model, while also contributing to better overall solution quality. Furthermore, although the constraint-gated loss does not yield significant improvement on GT_GAP, its positive effect on CONS_IF remains empirically valuable. These findings validate the effectiveness of the constraint-aware module and provide an explorative direction for enhancing the model’s sensitivity to constraint functions within problem embeddings.
TABLE III: Efficiency comparison between PAD, the LLM-based solver OPRO [38], and Gurobi [59]. The “Single-sample Inference Time” represents the lower and upper bounds of the latency observed for each method across all problems and input dimensions. “-” indicates metrics that cannot be precisely obtained or are not applicable to the corresponding method.
Property PAD ($p_{num}=2$) OPRO (GPT-5) Gurobi
GT_GAP $\downarrow$ $083\sim 0.2229 0.000$$\sim 0.8276 0.000$00.0000
CONS_IF $\uparrow$ $0.85\sim 1.00$ $0.50\sim 1.00$ $1.00$
Training /Building Time 2.78h Depends on manual coding